PROBLEM №5
A Japanese importer is planning to pay 1.2 mln British pounds for goods to be received. The current exchange rate is 1 GBP = 139.577 JPY or 1 JPY = 0.00716451 GBP
Please indicate what strategy with a use of options should be followed by a Japanese importer in order to have a protection against an adverse movement in the exchange rate.
In your analysis, you need to indicate what are the base and pricing currency.
For a chosen by you strike price (you need to justify the strike selected), estimate value of an option (type of an option, call of put, should be selected by you and justified), given that:
The annualized UK risk-free rate is r= 3.5 %, and the Japanese rate is r= 4%.
The time to expiration (T) is 0.25 years,
Historical volatility of the exchange rate σ=10%
Implied l volatility of the exchange rate σ=8.5%
Answer #1

Details provided : Amount to be paid = 1.2 mln pounds, current exchange rate : 1GBP = 139.577, UK risk-free rate = 3.5% Japanese risk-free rate = 4%, historical volatility of exchange rate = 10% and time = 0.25

a = e^(r-q)t = 1.001251

u = e^(volatility*squareroot(t)) = 1.051271

d = 1/u = 0.951229

p = (a-d)/(u-d) = 0.50005

current price of call option = 3.573

Since interest rate is higher in Japan it will depreciate yen against pounds and hence the importer should freeze the price with call option